OpenStudy (anonymous):

WILL MEDAL Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x^2 + 3 = -4x

OpenStudy (anonymous):

Out the '-4x' on the left to get a quadratic equation. You will get: \[x^2 + 4x + 3 = 0\] Now we can factorize to find solution. Factorize 3 to its components, you will get 1 and 3. And this 3 and 1 also add up to 4 (which is the coefficent of '4x'). So we can write the equation like this: \[(x+3)(x+1) = 0\] \[x + 3 = 0 => x_{1} = -3\] \[x + 1 = 0 => x_{2} = -1\] So our solution is pair of (-3, -1).